OpenStudy (anonymous):

what is the missing coordinate (-3,2), (0,y) and the slope is -2

OpenStudy (anonymous):

The formula for the slope of a straight line is = change of y/change of x. With your two coordinates, slope = (y - 2)/(0 - (-3)) or, slope = (y - 2)/3 Since th slope is given to be - 2, solve the equation (y - 2)/3 = - 2, and cross multiply, you get y - 2 = - 6, or y = - 4.